OpenStudy (anonymous):

combine: 4/y^2-9 + 5/y+3

OpenStudy (anonymous):

4 + 5y -15 -------------- y^2 - 9

OpenStudy (anonymous):

5y -11 ------ y^2 - 9

OpenStudy (anonymous):

the lcd = y^2 - 9 which expands to (x+3)(x-3) = 4 + 5(y-3) -------- y^2 - 9 = 5y - 11 ------- y^2 - 9